# A Comunication Framework for Compositional Generation

### v1

@@ 2.1 Lewis reconstruction game
The Lewis reconstruction game (Rita et al., 2022) is a special case of the Lewis Signaling Game (LSG) (Lewis, 1969) framework, in which two agents cooperate to reconstruct an observed object. Specifically, a sender agent parameterized by $\varphi$ observes an object $x$ from the set of objects $\mathcal{X}$ and produces a message $m\in\mathcal{M}$, where $\mathcal{M}$ is the set of all posible strings formed by concatenating symbols drawn from a finite vocabulary $V$. A receiver agent parameterized by $\omega$ observes $m$ and produces a reconsruction $x^{\prime}$. Both models are optimized to minimize a reconstruction loss $\mathcal{L}_{R}(x,x^{\prime})$.
@@
2.2 Iterated Learning
The IL model of cultural evolution (Kirby, 2001) proposes to emulate the emergence of language through the interactions of adult and new learning agents. In the original formulation, a space of pairs of signal + meaning is randomly generated, and in each iteration, a learning agent is partially exposed to the language learned by the adult agent in the previous iteration. The learning agent becomes the adult agent for the next iteration, and transmits a subset of its learned ”language”. By iterating this process, the initial random language gains structure and observable properties, such as compositionality.
This framework has been widely extended to include an additional task (Li & Bowling, 2019a; Ren et al., 2023; Zhou et al., 2022), instead of simply reshaping the language. Under this formulation, the training regime can be separated into two parts: Interaction (Int), where an agent is optimized to solve the additional task, and Imitation (Im), where a newly initialized agent ”learns” from the previous agent through a shared learned language.
2.3 Compositional data generation assumptions
In a formalism similar to (Ren et al., 2023), we first model the images by a random variable $X=\mathsf{GenX}(\mathbf{G},\mathbf{O})$, in which data $X$ are created via deterministic function $\mathsf{GenX}$ from a set of generating factors $\mathbf{G}$, and a source of randomness $\mathbf{O}$. We have access to samples from $X$ but not to $\mathsf{GenX},\mathbf{G},\mathbf{O}$, and the goal is to reproduce the distribution of $X$, i.e., generate a random variable $\widehat{X}$ that is close to $X$ in a natural metric $\mathcal{L}(X,\widehat{X})$.
We restrict to the case that $\mathbf{G}$ has a compositional structure, e.g. $\mathbf{G}=[G_{1},\dots,G_{n}]$ in which $G_{i}$ represent independent characteristics having finitely many possible values (as in the dataset of Sec. 5.1 for example). In this case, an efficient way to reproduce $X$ is to impose that $\widehat{X}$ is similarly generated as $\widehat{X}=\mathsf{GenX^{\prime}}(\mathbf{G^{\prime}},\epsilon)$ using a source of randomness $\epsilon$, where generating factors $\mathbf{G^{\prime}}$ should be a learned efficient encoding of $\mathbf{G}$. In fact we may argue, similar to (Ren et al., 2023), that if $\mathbf{G^{\prime}}$ is a compressed representation, then it tends to have low Kolmogorov complexity and to be isomorphic to the actual $\mathbf{G}$, leading to compositional generation of $\widehat{X}$. The arrow $X\to\mathbf{G^{\prime}}$ in the diagram below indicates the fact that $X$ is the datum from which elements from $\mathbf{G^{\prime}}$ have to be deduced for generating $\widehat{X}$, since $X$ is all we have access to.
$\mathbf{G}$$\mathbf{O}$$X$$\mathbf{G^{\prime}}$$\epsilon$$\widehat{X}$$\approx$ iso$\mathcal{L}(X,\widehat{X})$unknown
Here we focus on the case that $\mathbf{G}$ has a very large number of classes, in which case the main novel difficulty is to build a framework for finding the optimized compositional encoding $\mathbf{G^{\prime}}$ from observation of $X$ only. In order to avoid technical complications and confounding factors for our conclusions, we restrict to the simplified case that $\mathsf{GenX}$ generates a single element for each class in $\mathbf{G}$, in which case we can limit ourselves to the case of trivial noise sources $\mathbf{O},\epsilon$, and leave the challenges towards extension of the framework to cases of general $\mathbf{O},\epsilon$ to future study. Thus we assume that
$X=\mathsf{GenX}(\mathbf{G}),\quad\widehat{X}=\mathsf{GenX^{\prime}}(\mathbf{G^{\prime}}),$ (1)
@@ 2.3 Compositional data generation assumptions
In (1) the randomness of generated $X$ disappears, and effectively we work with a deterministic dataset $\mathcal{D}=\mathsf{Im}(\mathsf{GenX})\subseteq\mathcal{X}$ (where $\mathcal{X}$ is an ambient space, e.g. a ball in $\mathbb{R}^{D}$ for some large $D$) with one element for each class in $\mathbf{G}$, which we partition into $\mathcal{D}=\mathcal{D}_{train}\cup\mathcal{D}_{test}$ for training and testing. The compositional nature of $\mathbf{G}$ facilitates reconstruction up to isomorphism from just $\mathcal{D}_{train}$, a fact that is not guaranteed under general encodings (see §A for a formal treatment and Section 3.4 for discussion).
3 Proposed architecture
Figure 2: Overview of our proposed architecture: (Int) Sender $S_{\phi^{t}}$ and receiver $R_{\omega^{t}}$ are jointly trained to minimize the reconstruction error between the true state $x$ and predictions $\{x^{\prime}\}$ by encoding $x$ as a message $m^{t}_{x}$. (Im) The next sender $S_{\phi^{t+1}}$ is initialized to find a protocol that matches the final predicted state of $R_{\omega^{t}}(m^{t}_{x})$ and maximizes the pairwise distance between messages.
3.1 The signaling game approach
Our search space for $\mathbf{G^{\prime}}$ will consist of messages of length $\leq C$ consisting of words from a vocabulary $V$, i.e., we assume that $\mathbf{G^{\prime}}\in\mathcal{M}:=V^{C}$. The objective of finding the optimal encodings $\mathbf{G^{\prime}}$ is modeled as a signaling game: the map $S:\mathcal{X}\to\mathcal{M}$ plays the role of a sender that for each signal $x\in\mathcal{X}$ will produce some message $m\in V^{C}$, and $\mathsf{GenX^{\prime}}$ will be the restriction to $S(\mathcal{D})\subseteq\mathcal{M}$ of a suitable map $R:\mathcal{M}\to\mathcal{X}$, which represents a receiver that decodes the sender’s messages. The following diagram summarizes our notations:
Sender and receiver will be parametrized by recurrent neural networks, and have the common goal to find optimal signaling strategy, as measured by $\mathcal{L}(\cdot,\cdot)$.
In the context of the signaling game, the sender $S$ and receiver $R$ must establish a protocol that satisfies the following properties:
  1. (UE)
    Unambiguous Encoding: $S$ must produce an encoding of $x\in\mathcal{D}$ as messages $m\in\mathcal{M}$ that is unambiguous, i.e., injective. Specifically, states generated from distinct combinations of factors in $\mathbf{G}$ must be mapped to distinct messages by $S$.
  2. (EE)
    Efficient Encoding: The encoding of $m=S(x)$ should be such that $R$ is able to reconstruct any state $x\in\mathcal{D}$ by interpreting as few symbols from the message $m$ as possible. This property promotes efficiency in communication.
3.2 Unambiguity
As proposed by (Ren et al., 2023), compositionality emerges when $\mathbf{G}$ and $\mathbf{G^{\prime}}$ are isomorphic. In particular the map $\mathbf{G}\mapsto x\mapsto\mathbf{G^{\prime}}$ has to be injective. We assume the mapping $\mathbf{G}\mapsto x$ to be injective, thus the above composition being injective is equivalent to $x\mapsto\mathbf{G^{\prime}}$ being injective, which is condition UE.
However, achieving unambiguity is non-trivial. LSGs are known to exhibit two types of equilibria:
  • Separating Equilibria, where $S$ and $R$ coordinate to encode and decode messages that perfectly differentiate every state (UE).
  • Pooling Equilibria, where $S$ or $R$ associates multiple states with the same message.
Without appropriate incentive mechanisms, LSGs tend to converge to pooling equilibria due to their prevalence over separating equilibria (Huttegger et al., 2010b), a problem which (as shown in §D in an illustrative example) is exacerbated unless $|\mathcal{D}_{train}|$ remains relatively large.
This challenge is exacerbated in continuous state spaces. For instance, if two state vectors $x_{1}=\mathsf{GenX}(\mathbf{G}_{1}),$ and $x_{2}=\mathsf{GenX}(\mathbf{G}_{2})$, generated from $\mathbf{G}_{1},\mathbf{G}_{2}\in\mathcal{G}$, are sufficiently close (i.e., $d(x_{1},x_{2})<\delta$ for small $\delta$), typical loss functions give little incentive for $S$ to separate these states in $\mathcal{M}$: the best-case loss improvement that can occur when passing from coding $x_{1},x_{2}$ as the same message, to coding them as different messages, is bounded by $\sup_{x\in\mathcal{X}}|\mathcal{L}(x_{1},x)-\mathcal{L}(x_{2},x)|$, a value which tends to zero as $\delta\to 0$ in most situations of interest (e.g. whenever $\mathcal{L}$ is continuous and $\mathcal{D}\subset\mathcal{X}$ is bounded).
3.3 Efficiency
While EE is not strictly necessary for the theoretical soundness of the framework, it is a practical way to encourage generalization. This aligns with the information bottleneck principle and feature compression (Tishby et al., 2000; Kawaguchi et al., 2023). Additionally, the Kolmogorov complexity of the mapping $\mathbf{G^{\prime}}\mapsto\hat{X}$ depends on both the complexity of the model and the complexity of the inputs. Simplifying the message space reduces the complexity of the decoding scheme, facilitating learning and generalization.
3.4 Avoiding Holistic Encodings
A holistic encoding is an injective mapping $\mathcal{G}\to\mathcal{M}$ that encodes $\mathbf{G}$ into $\mathbf{G^{\prime}}$ but disregards the underlying characteristics of the elements (Kirby & Hurford, 2002) (see §A for a precise definition). For example, a holistic encoding might injectively enumerate each element $\mathbf{G}\in\mathcal{G}$ assigning to it a natural number $n_{\mathbf{G}}\in\mathbb{N}$. Since these mappings fail to reflect the structure of $\mathbf{G}$, they will fail to generalize well to unseen combinations of generating factors. Thus it is important to set up incentives for our communication game so that besides UE and EE it also achieves the following condition, which is favorable for compositional generalization:
  1. (NHE)
    Non Holistic Encodings: $S$ must avoid creating encodings that are holistic.
We assume that most combinations of the underlying factors are represented in some element of the training set $\mathcal{D}_{train}\subset\mathcal{D}$. This is reasonable to assume, even while assuming that $|\mathcal{D}_{train}|/|\mathcal{D}|$ is small, i.e., only a small portion of possible factor combinations are present (see §A.1 and in particular Thm. A.3 for a proof). In this case, compositional encodings could be deduced from $\mathcal{D}_{train}$ only. However, a naive optimization of efficiency objective EE on $\mathcal{D}_{train}$ necessarily leads to holistic encodings, a phenomenon that we illustrate in Example 3.1. Furthermore, in the large data $|\mathcal{D}|\gg 1$ regime holistic encodings have overwhelmingly higher Kolmogorov complexity than compositional ones, leading to poor generalization, we cannot guarantee UE and EE to generalize to $\mathcal{D}$. This is why enforcing NHE becomes necessary (for proofs of these claims and a formal discussion, see §B).
Example 3.1.
Consider a scenario with $k$ binary generating factors and a training set containing $|\mathcal{D}_{train}|=\frac{2^{k}}{8}=|\mathcal{D}|/8$ random samples from $\mathcal{D}$. The minimum number of bits required for a compositional mapping $\mathcal{G}\to\mathcal{M}$ is $\log 2^{k}=k$, whereas the minimum for a holistic mapping is $\log\frac{2^{k}}{8}=k-3$. Here, achieving EE solely over the training set is guaranteed to produce a holistic encoding that fails to generalize.
We then hypothesize that in order to favor NHE an effective heuristic is to maximize diversity (allowing larger ”useful lengths”) of encoded messages $\mathbf{G^{\prime}}$ (see §B).
4 Proposed architecture
Given a data distribution $\mathcal{X}$, a finite set of symbols $V$ and a fixed channel capacity $C$, we will define $\mathcal{M}=[V]^{C}$ as the set of all posible combinations of symbols in $V$ of length $C$, and sender and receiver $S$ and $R$ as learneable neural networks with parameters $\phi$ and $\omega$, respectively.
Following the IL framework (Kirby, 2001), we separate learning between an interaction phase and an imitation phase.
4.1 Interaction (Int)
During the Int phase, we jointly train $S$ and $R$ to optimize the reconstruction of $X$. To add compression pressure and enforce hierarchical representations and EE (Latheef et al., 2024), we include the following achitecture choices (i) we implement $S,R$ as recurrent neural networks, to further enforce the dependence of each symbol on the preceding ones; (ii) rather than optimizing the protocol to minimize the difference between the true state and the prediction given by the receiver after viewing the full message (Rita et al., 2022), the receiver makes a prediction with each new symbol from the message, and we minimize the average difference between the predicted states and the true states. For more explicit compression presure, we add a hyperparameter $\lambda\in[0,1]$. The Int phase loss becomes:
$\mathcal{L}^{int}_{\omega,\phi}:=\mathbb{E}_{x\sim\mathcal{X}}\left[\frac{1}{C}\sum_{i}^{C}\lambda^{i}d_{\mathcal{X}}(x,R_{\omega^{t}}(S_{\phi^{t}}(x)_{[i]}))\right]$ (2)
Where $[i]$ refer to the first $i$ symbols of a message and $d_{\mathcal{X}}$ will be a suitable distance or discrepancy measure over $\mathcal{X}$.
Remark 4.1 ($\mathcal{L}^{int}_{\omega,\phi}$ is a natural loss for compositional generative tasks).
If $x,x^{\prime}$ represent latent vectors of an accurate autoencoder (that maps related generators $\mathbf{G}$ to close-by vectors), then distance-based losses are natural for comparing $x,x^{\prime}$, as they are sensitive to positional information in latent space. This is not true for, e.g., log-likelihood losses, which are invariant under all measure-preserving transformations. It turns out that if we take log-likelihood-based signaling games loss as formalized, e.g., in (Rita et al., 2022) and we switch to the Wasserstein-type version where the expectation of $|x-x^{\prime}|^{p}$ replaces KL-discrepancy of the distributions of $x,x^{\prime}$ then we directly obtain the basic form of our loss $\mathcal{L}^{int}_{\omega,\phi}$ (see §C for the full derivation). In this sense, the form of the loss (2) is natural for our setting.
4.2 Imitation (Im)
During the Im phase, we apply a teacher-student regime to initialize the next iteration sender $S_{\phi^{t+1}}$. Specifically, frozen $S^{*}_{\phi^{t}}$ and $R^{*}_{\omega^{t}}$ generate ground truth predictions, which $S_{\phi^{t+1}}$ is trained to imitate, in a supervised setting.
Pooling (or mode collapse in the field of generative models) is a common problem in LSG in general and corresponds to the breaking of UE in optimization. As already mentioned in §3.2, this is exacerbated in pretrained generation encoders that map to a continuous space with weak guarantees of separation of the encodings. If there is not enough incentive for the sender to generate separate messages for similar states, then pooling is further favored by the compression pressure applied by our interaction scheme Int, in cases in which a strictly simpler encoding protocol can achieve marginally worse expected reconstruction.
UE requires us to enforce LSG separating equilibria, i.e., encodings that unambiguously represent every possible combination of generating factors. To push the iteration dynamics towards such equilibria, we employ the following three strategies that favor message diversity in the imitated protocol message space:
  1. 1.
    State space imitation: As opposed to other IL implementations (Ren et al., 2020; Li & Bowling, 2019a), the imitation accuracy is not tested in the message space, but rather in the state space, using the receiver network from the interaction phase. This allows more degrees of freedom for the imitation protocol.
  2. 2.
    Final state reconstruction: During the Im phase we use the following reconstruction loss:
    $\mathcal{L}^{rec}_{\phi^{t+1}}:=\mathbb{E}_{x\sim\mathcal{X}}\left[d_{\mathcal{X}}(x^{\star}_{t},x^{\prime}_{t+1})\right],$ (3)
[…]
    Differing from the Int phase, $\mathcal{L}^{rec}_{\phi^{t+1}}$ only includes the final predicted state, associated to the full message of length $C$. Along with the previous point, this should allow the student sender $S_{\phi^{t+1}}$ to more easily find different strategies with the same end result, ideally breaking pooling states encountered during the Int phase.
  3. 3.
    Pairwise distance maximization (PDM): To further favor diversity in the new protocol and encourage exploration, the student maximizes the Hamming distance between all messages in a given batch, pushing the sender to find the most ”diverse” protocol which behaves similarly to the teacher protocol, modulated by a hyperparameter $\beta$. With this, the Im loss becomes:
    $\mathcal{L}^{rec}_{\phi^{t+1}}+\beta\frac{1}{N_{batch}^{2}}\sum_{i,j}D[i,j],$ (4)
    $\text{where }D[i,j]:=d_{H}(S_{\phi^{t+1}}(x_{i}),S_{\phi^{t+1}}(x_{j})).$
    In §B.4 we prove that this simple regularization still gives a lower bound for entropy maximization (Williams & Peng, 1991), and an upper bound on the contrastive learning loss NT-Xtent (Chen et al., 2020).
5 Experiments
5.1 Datasets
We tested our framework using the composition split (Schott et al., 2022) of the Shapes3D (Matthey et al., 2017) dataset, which consists of colored images of 3D geometric shapes, with 6 underlying generating factors $G$: floor hue, wall hue, object hue, shape, scale and orientation. The total amount of atribute-value combinations is $480,000$.
The compositional split ensures that every possible value for each atribute is seen in the training set, yet most combinations of atribute values are reserved for the test set (See §A for a proof that this should hold with high probability).
5.2 Architecture
For the sender and receiver networks, we used the EGG Python library (Kharitonov et al., 2021). Both are single layer LSTMs (Hochreiter & Schmidhuber, 1997b) with an embedding size of 64 and hidden size of 256, followed by a two layer MLP with ReLU activation. We defined a vocabulary size of $|V|=25$, and a fixed channel length of $C=10$. For the channel we used a Gumbel-Softmax bottleneck (Jang et al., 2017) without straight-through estimation.
As a vision backbone, we pretrained a convolutional VAE for 2000 epochs with the Adam optimizer and a learning rate of $10^{-4}$. The encoder uses 4 convolutional layers with kernels in increasing powers of 2, from 32 to 256, with ReLU activation, and a final fully connected layer that outputs the predicted mean and logvariance vectors of dimension 256. The decoder is a mirrored version of the encoder with transposed convolutions, ReLU activation in the hidden layers, and sigmoid activation in the output layer.
Rather than training on the images themselves, we used the pretrained backbone to extract the predicted distributions for the latent vectors, and used these as our states. Since the states are multivariate Gaussians and not deterministic vectors, we take the reconstruction loss over the defined distributions. As mentioned in Rmk. 4.1, taking a latent space distance such as Wasserstein distance ($W_{2}$ in our case) is the most natural approach.
5.3 Training
The interaction-imitation games were run for 500 iterations, alternating one full epoch in each phase per iteration. We found that Int required smaller batch sizes to converge, wheras the Im phase PDM becomes a better entropy approximation at higher batch sizes, however scaling quadratically in computation. Therefore, we set the Int and Im batch sizes to 256 and 512, respectively.
To implement Im we defined a receiver $R_{\omega}$ and two senders $S_{\phi_{1}}$ and $S_{\phi_{2}}$ with their own optimizers. After participating in an Int iteration, $\phi^{1}$ and $\omega$ are frozen and used to train $S_{\phi_{2}}$ as described in §4.2. After Im, the weights $\phi_{2}$ are copied to $\phi_{1}$, and $\phi_{2}$ is reset. Importantly, the optimizers for $\phi_{1}$ and $\phi_{2}$ are never reset nor copied between them.
We compare our model with the following training schemes:
  • Interaction phase only (IPO): Training is performed without resetting the sender between iterations.
  • Interaction phase reset (IPR): After the Int phase, the sender is fully reset and not retrained with Im.
  • IL complete imitation (ILCI): Trained with IL framework using the full predicted sequence in the Im phase.
  • IL final state imitation (ILFI): Trained with IL framework using only the final state of the predicted sequence in the Im phase.
5.4 Metrics
Following (Chaabouni et al., 2022), we utilize the Topographic Similarity (TopSim) (Brighton & Kirby, 2006) between the emergent language and the ground truth generating factors, using Hamming distance for both. Furthermore, we measure the mean Hamming distance over all messages to detect pooling.
To quantify EE, we define the useful length $\ell_{\epsilon}(x)$ as the minimum message length for a given state $x$ and protocol encoded by $S_{\phi},R_{\omega}$, such that the additional accuracy gained by extending the message length by 1 is less than some small $\epsilon$, formally:
${\ell}_{\epsilon}(x):=\min\{i:d_{\mathcal{X}}(R_{\omega}(S_{\phi}(x)_{[i]}),R_{\omega}(S_{\phi}(x)_{[i+1]}))\leq\epsilon\},$
in which we restrict $i$ to $1,\dots,C$ and we take $\ell_{\epsilon}(x)=C$ if the minimum is not achieved in this range. Finally, we use the perceptual distance (Zhang et al., 2018) between generated images and the $W_{2}$ distance on the final predicted vector to measure reconstruction quality.
Figure 3: Comparison of validation reconstruction errors and evolution of useful length $\ell$ during training (averaged over 10 runs).
6 Results
Mode TopSim@200$\uparrow$ Percp. loss@200$\downarrow$
IPO 0.0182 0.4608
IPR 0.2385 0.2183
ILCI 0.2208 0.2126
ILFI 0.2339 0.1903
ILCI+PDM 0.2486 0.1944
ILFI+PDM (ours) 0.2665 0.1900
Table 1: Results from evaluating all tested methods with $\beta=0.5$ and $\lambda=0.95$ over the test set, averaged over 10 runs. To take into consideration that all methods had worsening results given enough iterations, we decided to take these measurements at epoch 200, where most models achieved their best performance.
6.1 Reconstruction performance
As evidenced in Fig. 3, our method achieves the lowest reconstruction error on the test set during early training, and becomes marginally suboptimal with respect to ILFI at the later stages of training. This behaviour is similar to the explore-exploit tradeoff of entropy maximization (Williams & Peng, 1991) in reinforcement learning.
Our method, however, achieves comparatively poor results within the training set, observed in Fig. 3 a), having a comparable full message-loss to IPR. Notably, tested methods without PDM achieve much better results during training, and slightly worse results in test. This behavior confirms our hypotheses about the need for additional inductive biases to enforce generalization in a compositional setting (see §B).
6.2 Impact of IL
As observed in Table 1 and Fig. 3, IL is required for convergence in this setting, as IPO resulted in near absolute pooling. This fact is evidenced by the useful length $\ell$ approaching zero, meaning most states are encoded into the same message.
Surprisingly, IPR consistently converged, which suggests that pooling equilibria are less stable to perturbations of the protocol, even when they are fully random. However, the lack of an Im phase results in lower quality image reconstructions and low TopSim similarity of the learned message space, as seen in Table 1.
6.3 ILCI vs ILFI
As observed in Table 1, ILFI obtains noticeably better results thtan ILCI, as measured by both reconstruction and TopSim. This result confirms our intuition of the importance of increasing the degrees of freedom order to maximize message expressiveness (see §B).
Fig. 3 shows that ILCI and ILFI have a much higher difference in reconstruction accuracy on the test set when not using DPM, behaving similarly on the train set regardless of regularization. This result suggests that utilizing ILFI has a higher impact on generalization than DPM.
6.4 Compression vs generalization
All tested methods converged to a smaller value of $\ell$ during training, with non PDM imitation methods closely approaching the theoretical minimum length $\lceil\log_{|V|}|\mathcal{D}|\rceil=5$ required to unambiguously encode every state. PDM appears to be damping this compression, avoiding full optimization of $\mathbb{E}[\ell]$. We claim this is required for NHE (see §B.3).
By tuning the parameters $\lambda$ and $\beta$, we were able to find a balance between compositionality and compression (See §2). For $\beta$ larger than $1$, the protocol does not converge efficiently, behaving in a similar manner to IPR. With $\lambda=0.95$ and $\beta=0.5$, we obtained the highest TopSim and the smallest perceptual loss on the training set among all tested methods (see Table 1).
7 Limitations and future work
As this work is the first proposal of an LSG + IL framework for generation tasks, several directions for future work emerged, and are outside the scope of this work:
  1. 1.
    For ease of rigorous evaluation, we performed a large ablation study on a toy dataset. In future work it is interesting to consider real-world data in the same framework. This will require an enriched (and plausibly valuable in itself) discussion of new ad-hoc evaluation metrics.
  2. 2.
    We considered a product-type compositional hypothesis, i.e., generating factors are assumed to be the outcome of independent components $\mathbf{G}=[G_{1},\dots,G_{n}]$. It is interesting to extend the theoretical framework (§A and 3.1) to more general compositional structures.
  3. 3.
    We assumed deterministically generated data as an injective function of $\mathcal{G}$. A clear avenue for future study lies in using a more general $\mathsf{GenX}$ as a randomized function with noise input as well as $\mathcal{G}$.
  4. 4.
    Generative tasks permit evaluations of novel compositional properties. We qualitatively explored the notable case message interpolation and found interpretable results (see §F.1 and §F.2). Exploring making quantitative such evaluations, seems a valuable avenue of future research in compositional generative machine learning.
8 Conclusion
In this work we presented a novel self-supervised framework for enforcing compositionality in generative models, applicable to probabilistic generative models such as VAEs.
By leveraging emergent language models, we created a tunable comunication bottleneck between collaborative agents, with ad-hoc biases towards compressibility and expressibility. Through this bottleneck, the latent vectors of probabilistic models could be compressed to a message space that shows good topological similarity with the hidden underlying generating space $\mathcal{G}$ of the dataset, indicating that the produced encodings inherited an approximate compositional structure.
We provided rigorous justifications and empirical evidence, in the form of a detailed ablation study, for the impact of our framework choices on the emergent properties of the comunication protocols.
Appendix A Recovering compositional structure from partially observed data
For the sake of clarity of treatment, we assume here that $\mathbf{G}=[G_{1},\dots,G_{n}]$ in which each factor $G_{i}$ has the same number $N$ of possible values, and as a consequence $\mathcal{G}=[1:N]^{n}$. Our results will extend directly to the general $\mathcal{G}$ case as shown in Rmk. A.4, but that statement seems too tedious for the reader to parse.

### v3

@@ A.2 Lewis reconstruction game
The Lewis reconstruction game (LRG) [47] is a special case of the Lewis Signaling Game (LSG) [33] framework, in which two agents cooperate to reconstruct an observed object. Specifically, a sender agent parameterized by $\varphi$ observes an object $x$ from an object space $\mathcal{X}$ and produces a message $m\in\mathcal{M}$, where $\mathcal{M}=V^{\star}$ is the set of all possible strings formed by concatenating symbols drawn from a finite vocabulary $V$. A receiver agent parameterized by $\omega$ observes $m$ and produces a reconstruction $x^{\prime}$. Both models are optimized to minimize a reconstruction loss $\mathcal{L}_{R}(x,x^{\prime})$.
@@
Appendix B Implementation Details & Hyperparameters
B.1 Baseline Model
  • Sender and Receiver: Both consist of a single-layer LSTM with an embedding size of 64 and a hidden size of 256. Outputs are passed through a two-layer MLP with ReLU activations.
  • Communication Channel: The vocabulary size is set to $|V|=15$ and messages are composed of $C=10$ discrete symbols. The Gumbel-Softmax bottleneck is used without straight-through estimation.
@@ B.1 Baseline Model
  • VAE Training: The VAE is pretrained for 15 epochs using the Adam optimizer with a learning rate of $1\times 10^{-3}$.
The interaction-imitation games were run for a maximum of 100 iterations, alternating one full epoch in each phase per iteration.
We used early stopping based on the validation MSE of the final reconstruction, starting from epoch $5$, with a minimum $\delta$ of $1\times 10^{-3}$ and patience of $5$. We used the Adam optimizer with default parameters and learning rates of $1\times 10^{-3}$ and $1\times 10^{-4}$ for Shapes3D [4] and MPI3D [17], respectively, as we found the training for the latter became unstable at higher learning rates.
We found that Interaction required smaller batch sizes to converge, whereas the Imitation phase PDM becomes a better entropy approximation at higher batch sizes, however scaling quadratically in computation. Therefore, we set the Interaction and Imitation batch sizes to 256 and 512, respectively.
To implement Imitation we defined a receiver $R_{\omega}$ and two senders $S_{\phi_{1}}$ and $S_{\phi_{2}}$ with their own optimizers. After participating in an Interaction iteration, $\phi^{1}$ and $\omega$ are frozen and used to train $S_{\phi_{2}}$ as described in §3.2. After Imitation, the weights $\phi_{2}$ are copied to $\phi_{1}$, and $\phi_{2}$ is reset. Importantly, the optimizers for $\phi_{1}$ and $\phi_{2}$ are never reset nor copied between them.
Appendix C Compute Resources
All experiments were conducted on our internal laboratory cluster, using GPUs with either 48 GB (NVIDIA A40) of VRAM. Each job was allocated a single GPU with 8 GB of VRAM usage on average, alongside access to a CPU with 20 cores, 118 GB of RAM, and local SSD storage for datasets and model checkpoints.
The experiments reported in this paper comprise 360 training runs, totaling approximately 180 GPU hours. Individual runs varied in length from 15 to 50 minutes, depending on early stopping criteria. These runs represent the finalized experiments whose results are presented in the main text.
In total, the research project required 2,310 experimental runs over the course of development, amounting to approximately 245 GPU days (5,880 GPU hours) of compute time. This includes preliminary experiments, ablations, and failed runs that were instrumental to model and protocol design but are not individually reported in the paper.
C.1 Useful length $\epsilon$
To define the useful length $\epsilon$ for both datasets, we observed reconstruction loss distribution over all positions in the message (see 3) for different $\lambda$ values, and estimated an average plateau point. For MPI3D we used a more conservative and precise threshold, as the reconstruction error varied more between methods.
Figure 3: Loss distribution over positions and $\lambda$ values for (a) Shapes3D and (b) MPI3D
Figure 4: Useful length values for different $\epsilon$ and $\lambda$ values for (a) Shapes3D and (b) MPI3D
Additionally, we analyzed the useful length graph for multiple $\epsilon$ values, and discarded those values where most models converged to either maximum or minimum possible length (see 4).
Appendix D Recovering compositional structure from partially observed data
For the sake of clarity of treatment, we assume here that $\mathbf{G}=[G_{1},\dots,G_{n}]$ in whichi factor $G_{i}$ has the same number $N$ of possible values, and as a consequence $\mathcal{G}=[1:N]^{n}$. Our results will extend directly to the general $\mathcal{G}$ case as shown in Rmk. D.4.